complete the square and find the two solutions. Simplfy the solutions and rationalize the denominators. x^2-3x-10=0
Answer by richwmiller(17219)

Don't need to complete the square.
Factor it.
(x-5)*(x+2)=0
x=5 and x=-2
complete the square.
x^2-3x-10=0
x^2-3x=10
(-3/2)^2
x^2-3x+9/4=49/4
(x-3/2)^2=49/4
x=3/2+7/2=10/2=5
x=3/2-7/2=-2
x=-2 and x=5